OpenStudy (anonymous):

Evaluate the difference quotient for the given function. Simplify your answer. 1/x using the formula f(x)-f(a)/x-a

OpenStudy (mathmale):

Dear Jasmine, welcome to OpenStudy. Please, would you provide the instructions for the problem you've posted? I can easily guess what the goal is, but would prefer that YOU state it for the benefit of yourself and those helping you.

OpenStudy (anonymous):

Evaluate the difference quotient for the given function. Simplify your answer.

OpenStudy (anonymous):

please help

OpenStudy (mathmale):

You want the difference quotient of f(x) = 1/x. OK. Clear. The difference quotient is defined by \[\frac{ f(x)-f(a) }{ x-a }\] where x is a variable (corresponding to points along the x axis) and a is a particular, constant x-value. f(x) = 1/x. What is f(a) = to?

OpenStudy (anonymous):

it doesn't give me fa

OpenStudy (mathmale):

You mean, "the problem statement doesn't give me f(a)." Read that as "function f(x) evaluated at x=a." f(x0 = 1/x here. Simply replace x with a on both sides of the equation. Result?

OpenStudy (anonymous):

1/a

OpenStudy (mathmale):

Right. Now write out (on paper) f(x) - f(a). Draw a line under that and underneath, write x-a. can you do that?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Now re-write the whole thing, replacing f(x) with 1/x and f(a) with 1/a.

OpenStudy (anonymous):

okay

OpenStudy (mathmale):

You now have \[\frac{ 1/x - 1/a }{ x-a }\] and must find and use the LCD of the numerator 1/x - 1/a to combine those two fractions through subtraction. Familiar with this step?

OpenStudy (mathmale):

Hint: 1/3 - 1/4 = (4-3)/12 = 1/12. Be certain you understand this procedure.

OpenStudy (anonymous):

no

OpenStudy (mathmale):

In 1/3 - 1/4 the two (different ) denominators are 3 and 4. Since those 2 have nothing in common, the LCD is 3*4=12, right?

OpenStudy (anonymous):

right

OpenStudy (mathmale):

So you need to do essentially the same thing with 1/x - 1/a. The 2 different denoms. are x and a; since they have nothing in common, the LCD here is xa. OK?

OpenStudy (mathmale):

Jasmine, I've checked and have found that you've apparently logged off OpenStudy. When you have someone helping you with a math problem and decide to log off, PLEASE be courteous and inform your helper.